Assume that 13​% of people are​ left-handed. Suppose 8 people are selected at random. Answer each question about​ right-handers below.

​e) What's the probability that the majority is​ right-handed?
Answer #1
here this is binomial with parameter n=8 and p=0.13

  probability that the majority is​right-handed =P(at most 3 are left handed):

probability = P(X<=3)= x=0x     (nCx)px(1−p)(n-x)    = 0.9871